# Universal Analog Computation: Fraïssé limits of dynamical systems

### v1

@@ Abstract.
Analog computation is an alternative to digital computation, that has recently re-gained prominence, since it includes neural networks. Further important examples are cellular automata and differential analyzers. While analog computers offer many advantages, they lack a notion of universality akin to universal digital computers. Since analog computers are best formalized as dynamical systems, we review scattered results on universal dynamical systems, identifying four senses of universality and connecting to coalgebra and domain theory. For nondeterministic systems, we construct a universal system as a Fraïssé limit. It not only is universal in many of the identified senses, it also is unique in additionally being homogeneous. For deterministic systems, a universal system cannot exist, but we provide a simple method for constructing subclasses of deterministic systems with a universal and homogeneous system. This way, we introduce sofic proshifts: those systems that are limits of sofic shifts. In fact, their universal and homogeneous system even is a limit of shifts of finite type and has the shadowing property.
@@
For inspiring discussions and comments, I am grateful to Herbert Jaeger, Michiel van Lambalgen, Marcus Pivato, Jan-Willem Romeijn, and the audiences at the Munich Center for Mathematical Philosophy (LMU Munich) and the at University of Groningen. Parts of the work were funded by the Bavarian Ministry for Digital Affairs.
@@ 1. Introduction
Analog computation is an alternative to digital computation, with a long and rich history [100, 14]. Examples include neural networks, the drivers of the present surge of artificial intelligence. They do not manipulate discrete symbols according to a bespoke program as in digital computation. Rather, they operate on real numbers according to their internal parameters, called weights, which typically are learned from data in a continuous updating process. Other famous examples are cellular automata, which distributively compute global behavior of cells via local update rules; or the differential analyzer, which computes by solving differential equations.
@@
In section 3, we motivate and define the notion of time-discrete, space-continuous, possibly nondeterministic dynamical system that we use. Specifically, a dynamical system here is a pair $(X,T)$ where $X$ is a zero-dimensional and compact metrizable space and $T:X\rightrightarrows X$ is a closed-valued and upper-hemicontinuos multifunction. We also call $(X,T)$ a nondeterministic system, and if $T$ is a function, we speak of a deterministic system. Section 4 describes the category of dynamical systems, whose morphisms are closely related to the notion of bisimulation. This notion of morphism is equivalent to that of a factor $f:(X,T)\to(Y,S)$, i.e., a surjective continuous function $f:X\to Y$ such that $y^{\prime}$ is an $S$-image of $y$ iff there are $x$ in $X$ and a $T$-image $x^{\prime}$ of $x$ such that $f(x)=y$ and $f(x^{\prime})=y^{\prime}$. For deterministic systems, this amounts to the usual notion of a factor, and this more general definition is also used for profinite graphs [57].
In section 5, we relate these systems to coalgebras and domain theory, thus contextualizing them within the two main general theories of computation. In particular, we show that our category of dynamical systems is equivalent to a category with a ‘domain-theoretic spirit’: the objects of this category are pairs $(A,\alpha)$ where $A$ is an algebraic lattice (with an additional property) and $\alpha:A\to A$ is a Scott-continuous function (with an additional property), and the morphisms $(\epsilon,\pi):(A,\alpha)\to(B,\beta)$ are embedding-projections pairs between $A$ and $B$ (with an additional property) such that $\pi\circ\beta\circ\epsilon\leq\alpha$ and $\alpha\circ\pi\leq\pi\circ\beta$. [^1] Note that these conditions are weaker than what is required for a coalgebra morphism between $(A,\alpha)$ and $(B,\beta)$, where one would require $\alpha\circ\pi=\pi\circ\beta$, which implies $\pi\circ\beta\circ\epsilon=\alpha\circ\pi\circ\epsilon=\alpha$.
Sections 68 establish the universality result for nondeterministic systems. To do so, we use the category-theoretic version of Fraïssé limits due to [32], which works in so-called algebroidal categories. In section 6, we state a theorem with sufficient conditions for algebroidality and conclude that the category of all dynamical systems is algebroidal. In section 7, we prove this algebroidality theorem. In section 8, we then use the Fraïssé limit to prove the following universality result:
Theorem (Corollary 8.4 below).
@@
    Universal: For any dynamical system $(Y,S)$, there is a factor $f:(U,T)\to(Y,S)$ (provided that $S$ is nontrivial, i.e., there is $y\in Y$ with $S(y)\neq\emptyset$).
@@
    Homogeneous: If $(Y,S)$ is a dynamical system over a finite set $Y$ and $f,f^{\prime}:(U,T)\to(Y,S)$ are two factors, then there is an isomorphism $\varphi:(U,T)\to(U,T)$ such that $f=f^{\prime}\circ\varphi$.
@@
However, the category of deterministic systems satisfies all other conditions for being algebroidal (i.e., it is semi-algebroidal). Thus, in section 10, we construct algebroidal subcategories. Any countable collection $S$ of shift spaces induces, by taking all limits along countable inverse sequences, an algebroidal category $\mathsf{DS}^{\mathsf{ef}}(S)$ of deterministic systems. A natural choice for $S$ is either the class $\mathsf{SFT}$ of shifts of finite type or the class $\mathsf{SOF}$ of sofic shifts. They are one of the two most studied classes of deterministic systems, and sofic shifts have close ties to automata theory. We call the systems in $\mathsf{DS}^{\mathsf{ef}}(\mathsf{SFT})$ (resp., $\mathsf{DS}^{\mathsf{ef}}(\mathsf{SOF})$) the $\omega$-proshifts of finite type (resp., the sofic $\omega$-proshifts). In section 11, we then show our second main universality result:
Theorem (Corollary 11.2 below).
@@
2. Four senses of universal analog computation
@@
Concerning 2, the differential analyzer—a revolutionary analog computer first built in 1931—was mathematically analyzed to precisely compute the solutions of algebraic differential equations [88, 83]. Later, [85] provides “an analogue, for analog computers, of the Universal Turing Machine” (p. 345), in the form of a concrete algebraic differential equation that is approximation universal in the sense that, for every continuous function $f$ on the real numbers and precision $\epsilon$, there is a smooth solution $y$ of that differential equation such that $y$ is $\epsilon$-close to $f$. Excitingly, [15] recently improved this to polynomial ordinary differential equations (which always have a unique and analytic solution). For neural networks, the universal approximation theorems [21, 53] similarly show that the functions realized by neural networks are dense in the set of all continuous functions.
@@
Concerning further related work, for a structurally similar discussion of universality of spin models, see [26], and for a category-theoretic framework for universality that includes spin models and Turing machines, see [44]. Moreover, our main tool, namely the Fraïssé-limit, has, beyond extensive use in model theory [52, ch. 7], also applications in the following fields: domain theory [32, 33], set theory [66], topos theory [17], graph theory [57, 16, 40], and dynamical systems theory [10, 69]—and see [73] for a more detailed overview. A stunning connection between these fields is provided by the KPT correspondence [62]: If $A$ is the Fraïssé-limit of a Fraïssé-order class $\mathcal{K}$, then $\mathcal{K}$ has a combinatorial property known as the Ramsey property iff the automorphism group of $A$ is extremely amenable, i.e., every continuous actions of this group on a compact Hausdorff space has a fixed point. This can also be used to construct universal flows for automorphism groups of Fraïssé limits. For a continuation of this line of research, see, e.g., [72, 75, 7, 76, 8, 63].
@@
Building on this rich body of work, our application of Fraïssé-limits is, to the best of our knowledge, new in constructing universal dynamical systems—motivated by the search for universal analog computation—in a way that unifies the case of nondeterministic systems and deterministic systems. Methodologically, we connect to the tools of theoretical computer science by using a category-theoretic version of the Fraïssé-limits and exploring links to domain theory and coalgebra. The closest existing results in the nondeterministic case is the work cited above on the projectively universal homogeneous (undirected) graph [57, 16, 40] (as an instance of the category-theoretic Fraïssé limit [32, 66]). Indeed, our nondeterministic systems can be seen as directed graphs. Although “some ideas [for undirected graphs] can be easily adapted to the more general case of directed graphs” [40, 3], we still include proofs, since the directed setting is conceptually central to dynamical systems and we reuse some results for the deterministic systems. The related work in the deterministic case is the literature cited above on KPT correspondence and genericity. But the focus is less on a universal system for actions of $\mathbb{N}$, but more on actions of automorphism groups of Fraïssé limits and understanding their universal minimal flow and their genericity. Moreover, we add category-theoretic tools to this investigation. For example, we corroborate the important role of shifts in dynamical systems theory by showing that they are the category-theoretically finite objects, and we define the category of, e.g., sofic $\omega$-proshifts as an instance of an important category-theoretic construction (essentially that of a pro-category). Thus, we add to a growing body of literature of using category theory for dynamical systems [70, 77, 9, 86].
3. Dynamical systems
The dynamical systems that we consider (to represent analog computers) are time-discrete, state-continuous, and possibly nondeterministic dynamical systems. In this section, we first state the formal definition, and then we motivate this choice and provide examples.
Nondeterminism means that a state might not have a unique next state, but that there is a set of potential next states. Hence the dynamics is a multifunction (or set-valued function). So we first recall the necessary definitions from set-valued analysis [3, ch. 17].
@@ 3. Dynamical systems
If $X$ and $Y$ are topological spaces, $F$ is closed-valued if each $F(x)$ is a closed subset of $Y$. The appropriate notion of continuity for us is that $F$ is upper-hemicontinuous if, for all $x\in X$ and open $V\subseteq Y$ with $F(x)\subseteq V$, there is an open $U\subseteq X$ with $x\in U$ and, for all $x^{\prime}\in U$, we have $F(x^{\prime})\subseteq V$. If $F$ is a function, it is continuous in the usual sense iff it is upper-hemicontinuous. The closed graph theorem [3, thm. 17.11] says: If $F:X\rightrightarrows Y$ is a multifunction between two compact Hausdorff spaces, then $F$ is upper-hemicontinuous and closed-valued iff the graph of $F$, i.e., the set $\mathrm{Gr}(F):=\{(x,y)\in X\times Y:y\in F(x)\}$, is closed. A corollary is that then the image $F[A]$ of a closed set $A\subseteq X$ is closed.
@@
To motivate this choice, topological dynamics is a well-established field studying dynamical systems as structures $(X,T)$ where $X$ is a compact Hausdorff space (usually metrizable) and $T:X\to X$ is a continuous function [101]. (This was pioneered by Poincaré studying differential equations also in the absence of analytic solutions.) Studying zero-dimensional topological systems is an important subfield [30]. This is not just because it is the setting of symbolic dynamics [71] but also because classical results like the Jewett–Krieger theorem show that the systems in ergodic theory (i.e., dynamics on probability spaces that preserve the measure) have zero-dimensional topological systems as counterparts. Here we generalize topological systems by also allowing nondeterministic dynamics. For simplicity, we still work with discrete time, i.e., $T$ describes the next states. Future work should generalize our results to continuous time (as used in continuous time analog computation) and to group or monoid actions more generally.
To illustrate our definition, let us briefly consider three examples.
@@
(1) A cellular automaton is a deterministic system $(\Sigma^{\mathbb{Z}^{n}},G)$ in the above sense. Here $\Sigma$ is the nonempty set of the finitely many states that a single cell can be in, so an element $c\in\Sigma^{\mathbb{Z}^{n}}$ describes the global configuration of the cellular automaton, i.e., the state of each cell in the $n$-dimensional grid $\mathbb{Z}^{n}$ of cells. The topology on $\Sigma^{\mathbb{Z}^{n}}$ is the product topology (where $\Sigma$ has the discrete topology). The function $G$ results from updating a global configuration according to the local updating rule. The local character renders $G$ continuous.
(2) The training dynamics $T$ of a neural network says how to update the current weights $w$ of the network, when presented with a batch of the training data, to a new choice of weights $w^{\prime}$. Formally, the dynamics is given by stochastic gradient descent, so the choice of batch is not deterministic, hence there are several possible next states $w^{\prime}$. Thus, the training dynamics $T:W\rightrightarrows W$ is a multifunction on the weight space.
@@
(3) Differential analyzers compute by solving differential equations and, as usual [98, sec. 6.2], such solutions are, if globally defined, dynamical systems $\Phi:\mathbb{R}\times X\to X$ with $X\subseteq\mathbb{R}^{n}$ open; or in discrete time steps, $T:=\Phi(1,\cdot):X\to X$. Practically, not much is lost when restricting to a closed and bounded subset of $X$ and deleting the countably many rational vectors: thus yielding a compact, zero-dimensional Polish space.
4. Categories of dynamical systems

### v2

@@ Abstract.
Analog computation is an alternative to digital computation, that has recently re-gained prominence, since it includes neural networks and neuromorphic computing. Further important examples are cellular automata and differential analyzers. While analog computers offer many advantages, they lack a notion of universality akin to universal digital computers. Since analog computers are best formalized as dynamical systems, we review scattered results on universal dynamical systems. We identify four senses of universality and connect to the two main general theories of computation: coalgebra and domain theory. For nondeterministic systems, we construct a universal system as a Fraïssé limit. It not only is universal in many of the identified senses, it also is unique in additionally being homogeneous. For deterministic systems, a universal system cannot exist, but we provide a simple method for constructing subclasses of deterministic systems with a universal and homogeneous system. This way, we introduce sofic proshifts: those systems that are limits of sofic shifts. In fact, their universal and homogeneous system even is a limit of shifts of finite type and has the shadowing property.
@@
For inspiring discussions and comments, I am grateful to Johan van Benthem, Herbert Jaeger, Michiel van Lambalgen, Hannes Leitgeb, Marcus Pivato, Jan-Willem Romeijn, and the audiences at the Munich Center for Mathematical Philosophy (LMU Munich) and the University of Groningen. Many thanks, in particular, to Marcus Pivato for very helpful comments on an earlier version of the manuscript. Parts of the work were funded by the Bavarian Ministry for Digital Affairs.
@@ 1. Introduction
Analog computation is an alternative to digital computation, with a long and rich history [116, 19]. Examples include neural networks, the drivers of the present surge of artificial intelligence. They do not manipulate discrete symbols according to a bespoke program as in digital computation. Rather, they operate on real numbers according to their internal parameters, called weights, which typically are learned from data in a continuous updating process. [^1] As an abstract model of computation, neural networks are analog. But when they are (physically) implemented, they usually are still simulated on digital hardware. However, developing analog or ‘neuromorphic’ hardware for neural networks is an active field of research [77]. More generally, neuromorphic computing develops computing systems which are inspired by the energy-efficient processing of biological brains. Other famous examples of analog computation are cellular automata, which distributively compute global behavior of cells via local update rules; or the differential analyzer, which computes by solving differential equations.
@@
In section 3, we motivate and define the notion of time-discrete, space-continuous, possibly nondeterministic dynamical system that we use. Specifically, a dynamical system here is a pair $(X,T)$ where $X$ is a zero-dimensional and compact metrizable space and $T:X\rightrightarrows X$ is a closed-valued and upper-hemicontinuos multifunction. We also call $(X,T)$ a nondeterministic system, and if $T$ is a function, we speak of a deterministic system. We introduce embedding-factor pairs as the morphisms between dynamical systems. Equivalently, a morphism is a factor $f:(X,T)\to(Y,S)$, i.e., a surjective continuous function $f:X\to Y$ such that $y^{\prime}$ is an $S$-image of $y$ iff there are $x$ in $X$ and a $T$-image $x^{\prime}$ of $x$ such that $f(x)=y$ and $f(x^{\prime})=y^{\prime}$ (cf. bisimulations). For deterministic systems, this amounts to the usual notion of a factor, and this more general definition is also used for profinite graphs [64]. With this framework in place, we formally state our main results in section 4.
In section 5, we relate our dynamical systems to coalgebras and domain theory, thus contextualizing them within the two main general theories of computation. In particular, we show that our category of dynamical systems is equivalent to a category with a ‘domain-theoretic spirit’: the objects of this category are pairs $(A,\alpha)$ where $A$ is an algebraic lattice (with an additional property) and $\alpha:A\to A$ is a Scott-continuous function (with an additional property), and the morphisms $(\epsilon,\pi):(A,\alpha)\to(B,\beta)$ are embedding-projections pairs between $A$ and $B$ (with an additional property) such that $\pi\circ\beta\circ\epsilon\leq\alpha$ and $\alpha\circ\pi\leq\pi\circ\beta$. [^2] Note that these conditions are weaker than what is required for a coalgebra morphism between $(A,\alpha)$ and $(B,\beta)$, where one would require $\alpha\circ\pi=\pi\circ\beta$, which implies $\pi\circ\beta\circ\epsilon=\alpha\circ\pi\circ\epsilon=\alpha$. In particular, to build universal systems, we cannot apply existence results of a final/terminal coalgebra [97, 2].
Sections 67 establish our universality result for nondeterministic systems. We use the category-theoretic version of Fraïssé limits due to [39], which works in so-called algebroidal categories. In section 6, prove a theorem with sufficient conditions for algebroidality and conclude that the category of all dynamical systems is algebroidal. In section 7, we then use the Fraïssé limit to prove the following universality result (generalizing one for undirected profinite graphs [64, 20, 47]):
Theorem (Theorem 4.2 below).
@@
    Universal: For any system $(Y,S)$, there is a factor $f:(U,T)\to(Y,S)$ (provided that $S$ is nontrivial, i.e., there is $y\in Y$ with $S(y)\neq\emptyset$).
@@
    Homogeneous: If $(Y,S)$ is a system over a finite set $Y$ and $f,f^{\prime}:(U,T)\to(Y,S)$ are two factors, then there is an isomorphism $\varphi:(U,T)\to(U,T)$ such that $f=f^{\prime}\circ\varphi$.
@@
However, the category of deterministic systems satisfies all other conditions for being algebroidal (i.e., it is semi-algebroidal). Thus, in section 9, we construct algebroidal subcategories. Any countable collection $S$ of shift spaces induces, by taking all limits along countable inverse sequences, an algebroidal category $\mathsf{DS}^{\mathsf{ef}}(S)$ of deterministic systems. A natural choice for $S$ is either the class $\mathsf{SFT}$ of shifts of finite type or the class $\mathsf{SOF}$ of sofic shifts. They are among the most studied classes of deterministic systems, and sofic shifts have close ties to automata theory. We call the systems in $\mathsf{DS}^{\mathsf{ef}}(\mathsf{SFT})$ (resp., $\mathsf{DS}^{\mathsf{ef}}(\mathsf{SOF})$) the $\omega$-proshifts of finite type (resp., the sofic $\omega$-proshifts). In section 10, we then show our second main universality result:
Theorem (Theorem 4.4 below).
@@
2. Background: Four senses of universal analog computation
@@
Concerning 2, the differential analyzer—a revolutionary analog computer first built in 1931—was mathematically analyzed to precisely compute the solutions of algebraic differential equations [101, 94]. Later, [96] provides “an analogue, for analog computers, of the Universal Turing Machine” (p. 345), in the form of a concrete algebraic differential equation that is approximation universal in the sense that, for every continuous function $f$ on the real numbers and precision $\epsilon$, there is a smooth solution $y$ of that differential equation such that $y$ is $\epsilon$-close to $f$. Excitingly, [18] recently improved this to polynomial ordinary differential equations (which always have a unique and analytic solution). For neural networks, the universal approximation theorems [26, 60] similarly show that the functions realized by neural networks are dense in the set of all continuous functions (see [17] for a recent result connecting to computability theory).
@@
Concerning further related work, for a structurally similar discussion of universality of spin models, see [32], and for a category-theoretic framework for universality that includes spin models and Turing machines, see [51]. Moreover, our main tool, namely the Fraïssé-limit, has, beyond extensive use in model theory [59, ch. 7], also applications in the following fields: domain theory [39, 40], set theory [75], topos theory [21], graph theory [64, 20, 47], and dynamical systems theory [13, 79]—and see [84] for a more detailed overview. A stunning connection between these fields is provided by the KPT correspondence [70]: If $A$ is the Fraïssé-limit of a Fraïssé-order class $\mathcal{K}$, then $\mathcal{K}$ has a combinatorial property known as the Ramsey property iff the automorphism group of $A$ is extremely amenable, i.e., every continuous action of this group on a compact Hausdorff space has a fixed point. This can also be used to construct universal flows for automorphism groups of Fraïssé limits. For a continuation of this line of research, see, e.g., [82, 86, 10, 87, 11, 73, 4].
@@
Building on this rich body of work, our application of Fraïssé-limits is, to the best of our knowledge, new in constructing universal dynamical systems—motivated by the search for universal analog computation—in a way that unifies the case of nondeterministic systems and deterministic systems. Methodologically, we connect to the tools of theoretical computer science by using a category-theoretic version of the Fraïssé-limits and exploring links to domain theory and coalgebra.
The closest existing results in the nondeterministic case is the work cited above on the projectively universal homogeneous (undirected) graph [64, 20, 47] (as an instance of the category-theoretic Fraïssé limit [39, 75]). Indeed, our nondeterministic systems can be seen as directed graphs. Although “some ideas [for undirected graphs] can be easily adapted to the more general case of directed graphs” [47, 3], we still include proofs, since the directed setting is conceptually central to dynamical systems and we reuse some results for the deterministic systems.
The related work in the deterministic case is the literature cited above on KPT correspondence and genericity. But the focus is less on a universal system for actions of $\mathbb{N}$, but more on actions of automorphism groups of Fraïssé limits and understanding their universal minimal flow and their genericity. Moreover, we add category-theoretic tools to this investigation. For example, we corroborate the important role of shifts in dynamical systems theory by showing that they are the category-theoretically finite objects, and we define the category of, e.g., sofic $\omega$-proshifts as an instance of an important category-theoretic construction (essentially that of a pro-category). Thus, we add to a growing body of literature of using category theory for dynamical systems [80, 88, 12, 99].
3. Categories of dynamical systems
We define the categories of dynamical systems that we will work with: subsection 3.1 defines the objects and subsection 3.2 the morphisms.
3.1. Objects: Dynamical systems
The dynamical systems that we consider (to represent analog computers) are time-discrete, state-continuous, and possibly nondeterministic dynamical systems. We first state the formal definition, and then we motivate this choice and provide examples.
Nondeterminism means that a state might not have a unique next state, but that there is a set of potential next states. Hence the dynamics is a multifunction (i.e., set-valued function). So we first recall the necessary definitions from set-valued analysis [5, ch. 17].
@@ 3.1. Objects: Dynamical systems
If $X$ and $Y$ are topological spaces, $F$ is closed-valued if each $F(x)$ is a closed subset of $Y$. The appropriate notion of continuity for us is that $F$ is upper-hemicontinuous if, for all $x\in X$ and open $V\subseteq Y$ with $F(x)\subseteq V$, there is an open $U\subseteq X$ with $x\in U$ and, for all $x^{\prime}\in U$, we have $F(x^{\prime})\subseteq V$. If $F$ is a function, it is continuous in the usual sense iff it is upper-hemicontinuous. The closed graph theorem [5, thm. 17.11] says: If $F:X\rightrightarrows Y$ is a multifunction between two compact Hausdorff spaces, then $F$ is upper-hemicontinuous and closed-valued iff the graph of $F$, i.e., the set $\mathrm{Gr}(F):=\{(x,y)\in X\times Y:y\in F(x)\}$, is closed. A corollary is that then the image $F[A]$ of a closed set $A\subseteq X$ is closed.
@@
To motivate this choice, topological dynamics is a well-established field studying dynamical systems as structures $(X,T)$ where $X$ is a Hausdorff space (typically compact and often also metrizable) and $T:X\to X$ is a continuous function [33]. (This was pioneered by Poincaré studying differential equations also in the absence of analytic solutions.) Studying zero-dimensional compact topological systems is an important subfield [36]. This is not just because it is the setting of symbolic dynamics [81] but also because classical results like the Jewett–Krieger theorem show that the systems in ergodic theory (i.e., dynamics on probability spaces that preserve the measure) have zero-dimensional topological systems as counterparts.
Moreover, as remark A.1 in appendix A.1 makes precise, we can also canonically turn the more general setting of a dynamics on a Polish space—like $\mathbb{R}^{n}$—into our setting. Specifically, if $X$ is a Polish space and $T:X\to X$ is a continuous function, then there is a—in a sense best possible—way to extend this into a deterministic system $(Y,S)$ in our sense. In particular, there is an injective function $\eta:X\to Y$ that is continuous with respect to a finer Polish topology on $X$ and which is equivariant, i.e., $\eta\circ T=S\circ\eta$.
Here we generalize topological systems by also allowing nondeterministic dynamics. For simplicity, we still work with discrete time, i.e., $T$ describes the next states. Future work should generalize our results to continuous time (as used in continuous time analog computation) and to group or monoid actions more generally.
To illustrate our definition, we consider, in the remainder of this section, some examples that highlight our computational perspective on dynamical systems.
Example 3.2 (Shifts).
Symbolic dynamics is an important subfield of dynamical systems theory. It developed from studying physical dynamical systems by discretizing both space and time, but it also found applications in storing and transmitting digital data [81]. The objects of study in symbolic dynamics are shifts, which are defined as follows. Let $A$ be a finite set, called alphabet. The (one-sided) full shift over $A$ is the system $(A^{\omega},\sigma)$ where $A^{\omega}$ carries the product topology of the discrete topology on $A$ and $\sigma$ is the shift map, i.e., $\sigma(x)_{n}=x_{n+1}$. A shift (aka shift space) is a closed, shift-invariant subset $X$ of a full shift on some alphabet. Hence, if $X$ is nonempty, $(X,\sigma)$ is in $\mathsf{det}\mathsf{Sys}^{\mathsf{ef}}$.
The two most studied classes of shifts are the following. The shifts of finite type are those shifts $X$ described by a finite list of forbidden words, i.e., $X$ is the set of those sequences $x$ over the alphabet $A$ of $X$ that do not contain a word $w$ from a finite set $F$ of words over $A$. Equivalently, shifts of finite type are those shifts whose sequences arise as the sequences of edges of infinite paths in a finite graph. The sofic shifts are those shifts that are factors of shifts of finite type. Equivalently, sofic shifts are those shifts whose sequences arise as the sequences of labels of infinite paths in a finite graph whose edges are labeled. This makes sofic shifts analogous to regular languages in automata theory.
Example 3.3 (Cellular automata).
@@
A cellular automaton is a deterministic system $(\Sigma^{\mathbb{Z}^{n}},G)$ in the above sense. Here $\Sigma$ is the nonempty set of the finitely many states that a single cell can be in, so an element $c\in\Sigma^{\mathbb{Z}^{n}}$ describes the global configuration of the cellular automaton, i.e., the state of each cell in the $n$-dimensional grid $\mathbb{Z}^{n}$ of cells. The topology on $\Sigma^{\mathbb{Z}^{n}}$ is the product topology (where $\Sigma$ has the discrete topology). The function $G$ results from updating a global configuration according to the local updating rule. The local character renders $G$ continuous.
For example, in dimension $n=1$ with binary states $\Sigma=\{0,1\}$, the well-known rule “110” says that a cell changes from 0 to $1$ precisely if the cell to its right is $1$ (“life spreads left’), and a cell changes from $1$ to 0 precisely if both the cell to the left and to the right is $1$ (“overpopulation”). [24] shows that this simple rule yields a Turing universal cellular automaton.
Example 3.4 (Differential equations).
@@
Differential analyzers compute by solving (ordinary) differential equations—coming, e.g., from engineering [19, 116]. As usual [114, sec. 6.2], such solutions are, if globally defined, dynamical systems $\Phi:\mathbb{R}\times X\to X$ with $X\subseteq\mathbb{R}^{n}$ open; or in discrete time steps, $T:=\Phi(1,\cdot):X\to X$. As described by remark A.1, we can view this dynamics on the Polish space $X$ in the larger, compact and zero-dimensional space $Y$, rendering it a deterministic system in our formal sense.
Example 3.5 (Training dynamics).
One may view the training process of a neural network as computing weights for the neural networks with which it solves its task well. The training dynamics $T$ takes the current weights $w$ of the network and a batch of the training data, and then produces (using backpropagation) updated weights $w^{\prime}$. Formally, the dynamics is given by stochastic gradient descent, so the choice of batch is not deterministic, hence there are several possible next states $w^{\prime}$. Thus, the training dynamics $T:W\rightrightarrows W$ is a multifunction on the weight space $W$. Understanding this dynamics is an important problem in the theory of machine learning [98, 66].
Example 3.6 (Reservoir computing).
Reservoir computing is one prominent approach to computing on non-digital hardware. The basic idea is that inputs are mapped into a high-dimensional state space of a nonlinear dynamical system—called the reservoir—and only a simple read-out function is trained to map from the high-dimensional state space to outputs [112]. This was initiated independently by [67] and [83], with a rich history since. For the purpose of this example, let us consider the first approach. [^6] In the case where there is no feedback from the output back to the reservoir. The reservoir is a recurrent neural network with weight matrix $W$ whose state, at time $t\in\{0,1,2,\ldots\}$, is the vector $x(t)$ describing the activation value of each neuron. Additionally, there are input neurons whose state, at time $t$, is the vector $u(t)$ describing the value of each input neuron. (So the input can change over time, i.e., is sequential.) A matrix $W^{\text{in}}$ maps input neuron states into states of the reservoir. The state of the reservoir at the next time step is computed by
(1) $x(t+1):=f\big(W^{\text{in}}u(t+1)+Wx(t)\big),$
where $f$ is the function that applies to each neuron its activation function (e.g., a sigmoid). Moreover, there are output neurons whose activations, at time $t$, are given by the vector $y(t)=W^{\text{out}}x(t)$. Only the weight matrix $W^{\text{out}}$ is updated during training. If the input sequence $u$ is not fixed, we can view this as a nondeterministic system $T:X\rightrightarrows X$ on the state space $X$ of the reservoir system, with $T(x)$ being the set of next states realized by some input vector as in 1.
3.2. Morphisms: Embedding-factor pairs